OpenStudy (anonymous):

What is the product of 6the square root of the quantity 12 times k. • 4kthe square root of 3.? Simplify if possible.

OpenStudy (anonymous):

help anyone

OpenStudy (phi):

\[ 5 \sqrt{12k}\cdot 4k\sqrt{3} \] everything is being multiplied, so you can change the order (because 2*3 is the same as 3*2, for example) \[ 5\cdot 4k \sqrt{12k}\cdot\sqrt{3} \] 5 times 4 is 20, so you can simplify this to \[ 20k \sqrt{12k}\cdot\sqrt{3} \] you could "merge" the square roots because \[ \sqrt{a\cdot b}= \sqrt{a} \cdot \sqrt{b} \] and vice versa so you could write this as \[ 20k \sqrt{3\cdot 12k} \] or \[ 20k \sqrt{36k} \] can you continue ?